# High-dimensional dynamical systems: co-existence of attractors, phase transitions, maximal Lyapunov exponent and response to periodic drive

### v1

@@ 1 Introduction
In the past forty years there has been growing interest in dynamical systems where the number of degrees of freedom is large. This high-dimensional regime is rather different from the one that is typically considered in dynamical system theory and it is relevant in the study of a large class of complex systems. The study of dynamical systems in high dimension was started in the foundational work by Sompolinsky, Crisanti and Sommers [2, 14]. This work considers simple models of random recurrent neural networks and shows that depending on the strength of the interactions between the neurons in the system, the asymptotic state of the dynamics can be either a fixed point attractor, where the degrees of freedom converge to a particular random point in phase space, or a chaotic attractor where endogenous motion is kept steady. Most importantly, this work shows how to use field theory techniques to study the properties of the chaotic attractor, in particular the Maximal Lyapunov Exponent (MLE). Field theory techniques to study neural networks are also the subject of recent reviews, see [15, 16]. The class of models studied in [2] became the building blocks of several subsequent studies of recurrent neural networks, including recent works [17, 18, 19, 20, martorell2024dynamically, clark2024theory, 21, 22].
@@ 2.1 A large class of models
In Eq. (2.4), we have assumed that $f$ admits a Taylor expansion with positive semidefinite Taylor coefficients. As soon as $f$ is non-linear, the corresponding dynamical system is non-linear. We call pure models the ones for which $f(z)$ is a pure monomial. Conversely, we address the corresponding models as mixed. This terminology is directly derived from spin glasses [28, 29] where the random fields $\boldsymbol{r}$} are actually the gradient of a spin glass Hamiltonian and Eq. (2.1) describes the gradient flow dynamics on the corresponding energy landscape. The control parameter $g$ sets the strength of the non-linear forcing of the dynamics. We assume, without losing generality, that $f$ does not contain a constant term. Indeed this case would be equivalent with the model with an additional external field. Furthermore, we assume that $f$ is not a linear function. This case corresponds to having a dynamical system that is integrable. We are interested in non-linear cases that cannot be understood in terms of an extensive number of uncoupled/weakly coupled degrees of freedom.
@@ 2.1.1 Initial condition of the dynamical systems
$\begin{split}\int_{-\infty}^{\infty}{\rm d}zp_{0}(z)z^{2}&=\mathtt{C}\\ \int_{-\infty}^{\infty}{\rm d}zp_{0}(z)z^{2}&=\mathtt{m}\>.\end{split}$ (2.6)
@@ 3.1 Derivation of the DMFT equations
Although there are many ways to derive the DMFT equations, here we follow a route typical in statistical physics of out of equilibrium systems and transform the dynamical equations in a path integral [32]. In the large $N$ limit, this path integral can be evaluated via a saddle point method. The corresponding saddle point equations are the DMFT equations.
@@ 3.1.1 Path integral formalism
${\mathbb{E}}\exp\left[i\int_{t}\hat{x}_{i}(t)r_{i}({\boldsymbol{x}}(t))\right]=\exp\left[-\frac{g^{2}}{2}\sum_{i}\int_{t,t^{\prime}}\hat{x}_{i}(t)f(C(t,t^{\prime}))\hat{x}_{i}(t^{\prime})\right]\>.$ (3.3)
@@
${\mathbb{E}}{Z}[\hat{{\boldsymbol{b}}},{\boldsymbol{b}}]=\int_{C(0,0)=\mathtt{C};m(0)=\mathtt{m}}\mathcal{D}C\mathcal{D}\hat{C}\mathcal{D}m\mathcal{D}\hat{m}\,e^{-N\mathcal{L}[{{\boldsymbol{x}}},{\hat{{\boldsymbol{x}}}},C,\hat{C},m,\hat{m},{\boldsymbol{b}},\hat{\boldsymbol{b}}]}$ (3.6)
@@
$\begin{split}\mathcal{L}[{{\boldsymbol{b}}},{\hat{{\boldsymbol{b}}}},C,\hat{C},m,\hat{m}]&=i\int_{t,t^{\prime}\geq 0}{\rm d}t{\rm d}t^{\prime}\hat{C}(t,t^{\prime})C(t,t^{\prime})+i\int_{t\geq 0}{\rm d}t\hat{m}(t)m(t)-W[{{\boldsymbol{b}}},{\hat{{\boldsymbol{b}}}},C,\hat{C},m,\hat{m}]\\ W[{{\boldsymbol{b}}},{\hat{{\boldsymbol{b}}}},C,\hat{C},m,\hat{m}]&=\frac{1}{N}\sum_{k}\ln\int\mathcal{D}x_{k}\mathcal{D}\hat{x}_{k}\,e^{-S_{k}[x_{k},\hat{x}_{k},C,\hat{C}]+\sum_{k=1}^{N}\int_{t\geq 0}{\rm d}t\left[i\hat{x}_{k}(t)b_{k}(t)+i\hat{b}_{k}(t)x_{k}(t)\right]}\\ &\equiv\frac{1}{N}\sum_{k=1}^{N}\ln\mathcal{Z}_{k}[{b_{k}},{\hat{b}_{k}},C,\hat{C},m,\hat{m}]\\ S_{k}[x_{k},\hat{x}_{k},C,\hat{C},m,\hat{m}]&=\int_{t\geq 0}{\rm d}t\,\left(i\hat{x}_{k}(t)\left[\partial_{t}x_{k}(t)+\hat{\mu}(C(t,t))x_{k}(t)-J_{0}m(t)-h\right]-i\hat{m}(t)x_{k}(t)\right)\\ &-\int_{t,t^{\prime}\geq 0}{\rm d}t{\rm d}t^{\prime}\left(-\frac{g^{2}}{2}\hat{x}_{k}(t)f(C(t,t^{\prime}))i\hat{x}_{k}(t^{\prime})+x_{k}(t)\hat{C}(t,t^{\prime})x_{k}(t^{\prime})\right)\>.\end{split}$ (3.7)
@@
where we have indicated with $\mathcal{Z}^{(0)}_{k}=\mathcal{Z}_{k}[0,0,C,\hat{C},m,\hat{m}$] evaluated at the saddle point. Causality of the dynamics imposes that $\langle i\hat{x}_{k}(t)i\hat{x}_{k}(t^{\prime})\rangle_{\mathcal{Z}_{k}^{(0)}}=0$ and $\langle i\hat{x}_{k}(t)x_{k}(t)\rangle_{\mathcal{Z}_{k}^{(0)}}=0$. This implies that $\hat{C}(t,t^{\prime})=0$. On the same lines, causality implies that $\langle x_{k}(t)\rangle_{\mathcal{Z}_{k}^{(0)}}=0$ and therefore $\hat{m}(t)=0$. Therefore the saddle point equations reduce to Eqs. (3.9) and 3.11.
@@ 3.1.2 Self-consistent stochastic process
where the noise $\eta(t)$ is a Langevin process with zero mean and variance given by
@@ 3.2 Discrete time dynamical system
$\begin{split}C(t+{\rm d}t,t^{\prime})&=C(t,t^{\prime})+{\rm d}t\left[-\mu(t)C(t,t^{\prime})+g^{2}{\rm d}t\sum_{n=0}f(C(t,n{\rm d}t))R(t^{\prime},n{\rm d}t)\right.\\ &\left.+(J_{0}m(t)+h)m(t^{\prime})\right]\ \ \ \ \ \ \ t>t^{\prime}\\ C(t+{\rm d}t,t+{\rm d}t)&=C(t,t)+2{\rm d}t\left[-\mu(t)C(t,t)+g^{2}{\rm d}t\sum_{n=0}f(C(t,n{\rm d}t))R(t,n{\rm d}t)\right]\\ &+2{\rm d}t(J_{0}m(t)+h)m(t)+{\rm d}t^{2}L(t)\\ R(t+{\rm d}t,t^{\prime})&=R(t,t^{\prime})-{\rm d}t\mu(t)R(t,t^{\prime})+\delta_{t,t^{\prime}}\\ \mu(t)&=\hat{\mu}(C(t,t))\\ m(t+{\rm d}t)&=m(t)+{\rm d}t\left[(J_{0}-\mu(t))m(t)+h\right]\\ L(t)&=\mu^{2}(t)C(t,t)+(J_{0}m(t)+h)^{2}-\mu(t)(J_{0}m(t)+h)m(t)\\ &-2g^{2}\mu(t)\sum_{n=0}f(C(t,n{\rm d}t))R(t,n{\rm d}t){+g^{2}f(C(t,t))}\end{split}$ (3.24)
@@ 6.1 DMFT with periodic drive
$\begin{split}x_{i}^{(0)}(t)&=-\mu(t)x_{i}^{(0)}(t)+H_{i}(t)\\ x_{i}^{(1)}(t)&=-\mu(t)x_{i}^{(1)}(t)+gr_{i}({\boldsymbol{x}}(t))\>.\end{split}$ (6.4)
@@
${\mathbb{E}}\left[\cos(\omega s+\theta)\cos(\omega s^{\prime}+\theta)\right]=\frac{1}{2}\cos(\omega(t-t^{\prime})),$ (6.9)
@@
$\begin{split}\partial_{t}C^{(1)}(t,t^{\prime})&=-\mu(t)C^{(1)}(t,t^{\prime})+g^{2}\int_{0}^{t}{\rm d}sf(C(t,s))R^{(1)}(t^{\prime},s)\\ \partial_{t}R^{(1)}(t,t^{\prime})&-\mu(t)R^{(1)}(t,t^{\prime})+\delta(t-t^{\prime})\>.\end{split}$ (6.13)
@@
This is because one can show that, in the $N\to\infty$ limit, ${\boldsymbol{x}}^{(0)}(t)\cdot{\boldsymbol{x}}^{(1)}(t^{\prime})/N=0$.
@@ 7 Maximum Lyapunov exponent
Chaotic attractors can be characterized by their maximal Lyapunov exponent (MLE), which describes how trajectories are sensible to small perturbations in their initial condition.
@@
In this section, we study the problem of determining the MLE for the class of models we are focusing on. We show a set of results:
  • We develop a formalism that is alternative to the one of [2] and that avoids the field theory techniques. Our techniques are similar to [15] but we are able to get a more precise description of the MLE and the divergence of the distance between two closely initialized trajectories.
  • The formalism allows to derive a set of equations for the MLE that are similar to the ones of [2].
  • For the simplest dynamical system, we can obtain the MLE exactly in the whole chaotic phase and characterize its behavior at phase transition points.
@@ 7.1 Definition of the maximal Lyapunov exponent
$\delta{\boldsymbol{x}}(t)\simeq{\varepsilon}({\boldsymbol{b}}_{1}-{\boldsymbol{b}}_{2})\,e^{\Lambda_{0}(t-t_{0})}\ \ \ \ t\gg t_{0}$ (7.2)
@@ 7.2 DMFT analysis
The distance $D(t,t_{0})$ is a self-averaging quantity since it is a sum of positive semidefinite terms, and therefore when $N\to\infty$ it concentrates on its mean. Thus, it can be computed via DMFT. The strategy that we use consists in implementing the DMFT analysis for two copies of the system. The two copies are initialized at the same initial point, perform the same trajectory up to time $t_{0}$, get perturbed by the addition of two infinitesimally small random vectors, and then are let free to evolve.
@@ 7.2.1 Two replicas formalism
$x_{\alpha}(t_{0}+{\rm d}t)=x_{\alpha}(t_{0})+{\rm d}t\left[-\mu_{\alpha}(t)x_{\alpha}(t)+\eta_{\alpha}(t)+J_{0}m_{\alpha}(t)+h\right]{\varepsilon}b_{\alpha},$ (7.10)
@@
$\begin{split}H_{\alpha}(t)&=\begin{cases}{\varepsilon}r_{\alpha}&t=t_{0}\\ 0&t\neq t_{0}\>.\end{cases}\end{split}$ (7.12)
@@
$\hat{H}_{\alpha}(t)={\varepsilon}r_{\alpha}\delta(t-t_{0})$ (7.13)
@@ 7.2.3 Symmetric two replicas DMFT equations
$C_{d}(t+{\rm d}t,t^{\prime})$ $=C_{d}(t,t^{\prime})+{\rm d}t\left[-\tilde{\mu}(t)C_{d}(t,t^{\prime})+g^{2}{\rm d}t\sum_{n=0}f(C_{d}(t,n{\rm d}t))R_{d}(t^{\prime},n{\rm d}t)\right$. (7.22)
@@
$C_{o}(t+{\rm d}t,t^{\prime})$ $=C_{o}(t,t^{\prime})+{\rm d}t\left[-\tilde{\mu}(t)C_{o}(t,t^{\prime})+g^{2}{\rm d}t\sum_{n=0}f(C_{o}(t,n{\rm d}t))R_{d}(t^{\prime},n{\rm d}t)\right$. (7.23)
@@
$C_{d}(t+{\rm d}t,t+{\rm d}t)$ $=C_{d}(t,t)+2{\rm d}t\left[-\tilde{\mu}(t)C_{d}(t,t)+g^{2}{\rm d}t\sum_{n=0}f(C_{d}(t,n{\rm d}t))R_{d}(t,n{\rm d}t)\right]$ (7.24)
@@
$C_{o}(t+{\rm d}t,t+{\rm d}t)$ $=C_{o}(t,t)+2{\rm d}t\left[-\tilde{\mu}(t)C_{o}(t,t)+g^{2}{\rm d}t\sum_{n=0}f(C_{o}(t,n{\rm d}t))R_{d}(t,n{\rm d}t)\right]$ (7.25)

### v2

# High-dimensional random dynamical systems: co-existence of chaotic attractors, phase transitions, and response to periodic drive

@@ 1 Introduction
In the past forty years there has been growing interest in dynamical systems where the number of degrees of freedom is large. This high-dimensional regime is rather different from the one that is typically considered in dynamical system theory and it is relevant in the study of a large class of complex systems. The study of dynamical systems in high dimension was started in the foundational work by Sompolinsky, Crisanti and Sommers [43, 4]. This work considers simple models of random recurrent neural networks and shows that depending on the strength of the interactions between the neurons in the system, the asymptotic state of the dynamics can be either a fixed point attractor, where the degrees of freedom converge to a particular random point in phase space, or a chaotic attractor where endogenous motion is kept steady. Most importantly, this work shows how to use field theory techniques to study the properties of the chaotic attractor, in particular the Maximal Lyapunov Exponent (MLE). Field theory techniques to study neural networks are also the subject of recent reviews, see [21, 41]. The class of models studied in [43] became the building blocks of several subsequent studies of recurrent neural networks, including recent works [28, 35, 2, 15, 27, 3, 22, 20].
@@ 2.1 A large class of models
In Eq. (2.4), we have assumed that $f$ admits a Taylor expansion with positive semidefinite Taylor coefficients. As soon as $f$ is non-linear, the corresponding dynamical system is non-linear. We call pure models the ones for which $f(z)$ is a pure monomial. Conversely, we address the corresponding models as mixed. This terminology is directly derived from spin glasses [32, 5, 1, 12] where the random fields $\boldsymbol{r}$} are actually the gradient of a spin glass Hamiltonian and Eq. (2.1) describes the gradient flow dynamics on the corresponding energy landscape. The control parameter $g$ sets the strength of the non-linear forcing of the dynamics. We assume, without losing generality, that $f$ does not contain a constant term. Indeed this case would be equivalent with the model with an additional external field. Furthermore, we assume that $f$ is not a linear function. This case corresponds to having a dynamical system that is integrable, see [6, 44]. We are interested in non-linear cases that cannot be understood in terms of an extensive number of uncoupled/weakly coupled degrees of freedom.
@@ 2.1.1 Initial condition of the dynamical systems
$\begin{split}\int_{-\infty}^{\infty}{\rm d}zp_{0}(z)z^{2}&=\mathtt{C}\\ \int_{-\infty}^{\infty}{\rm d}zp_{0}(z)z&=\mathtt{m}\>.\end{split}$ (2.6)
@@ 3.1 Derivation of the DMFT equations
Although there are many ways to derive the DMFT equations, here we follow a route typical in statistical physics of out of equilibrium systems and transform the dynamical equations in a path integral [26, 11, 9, 48]. In the large $N$ limit, this path integral can be evaluated via a saddle point method. The corresponding saddle point equations are the DMFT equations.
@@ 3.1.1 Path integral formalism
${\mathbb{E}}\exp\left[i\int_{t}\hat{x}_{i}(t)r_{i}({\boldsymbol{x}}(t))\right]=\exp\left[-\frac{g^{2}}{2}\int_{t,t^{\prime}}\hat{x}_{i}(t)f(C(t,t^{\prime}))\hat{x}_{i}(t^{\prime})\right]\>.$ (3.3)
@@
${\mathbb{E}}{Z}[\hat{{\boldsymbol{b}}},{\boldsymbol{b}}]=\int_{C(0,0)=\mathtt{C};m(0)=\mathtt{m}}\mathcal{D}C\mathcal{D}\hat{C}\mathcal{D}m\mathcal{D}\hat{m}\,e^{-N\mathcal{L}[{{\boldsymbol{b}}},{\hat{{\boldsymbol{b}}}},C,\hat{C},m,\hat{m}]}$ (3.6)
@@
$\begin{split}\mathcal{L}[{{\boldsymbol{b}}},{\hat{{\boldsymbol{b}}}},C,\hat{C},m,\hat{m}]&=i\int_{t,t^{\prime}\geq 0}{\rm d}t{\rm d}t^{\prime}\hat{C}(t,t^{\prime})C(t,t^{\prime})+i\int_{t\geq 0}{\rm d}t\hat{m}(t)m(t)-W[{{\boldsymbol{b}}},{\hat{{\boldsymbol{b}}}},C,\hat{C},m,\hat{m}]\\ W[{{\boldsymbol{b}}},{\hat{{\boldsymbol{b}}}},C,\hat{C},m,\hat{m}]&=\frac{1}{N}\sum_{k}\ln\int\mathcal{D}x_{k}\mathcal{D}\hat{x}_{k}\,e^{-S_{k}[x_{k},\hat{x}_{k},C,\hat{C}]+\int_{t\geq 0}{\rm d}t\left[i\hat{x}_{k}(t)b_{k}(t)+i\hat{b}_{k}(t)x_{k}(t)\right]}\\ &\equiv\frac{1}{N}\sum_{k=1}^{N}\ln\mathcal{Z}_{k}[{b_{k}},{\hat{b}_{k}},C,\hat{C},m,\hat{m}]\\ S_{k}[x_{k},\hat{x}_{k},C,\hat{C},m,\hat{m}]&=\int_{t\geq 0}{\rm d}t\,\left(i\hat{x}_{k}(t)\left[\partial_{t}x_{k}(t)+\hat{\mu}(C(t,t))x_{k}(t)-J_{0}m(t)-h\right]-i\hat{m}(t)x_{k}(t)\right)\\ &-\int_{t,t^{\prime}\geq 0}{\rm d}t{\rm d}t^{\prime}\left(-\frac{g^{2}}{2}\hat{x}_{k}(t)f(C(t,t^{\prime}))i\hat{x}_{k}(t^{\prime})+x_{k}(t)\hat{C}(t,t^{\prime})x_{k}(t^{\prime})\right)\>.\end{split}$ (3.7)
@@
where we have indicated with $\mathcal{Z}^{(0)}_{k}=\mathcal{Z}_{k}[0,0,C,\hat{C},m,\hat{m}$] evaluated at the saddle point. Causality of the dynamics imposes that $\langle i\hat{x}_{k}(t)i\hat{x}_{k}(t^{\prime})\rangle_{\mathcal{Z}_{k}^{(0)}}=0$ and $\langle i\hat{x}_{k}(t)x_{k}(t)\rangle_{\mathcal{Z}_{k}^{(0)}}=0$. This implies that $\hat{C}(t,t^{\prime})=0$. On the same lines, causality implies that $\langle\hat{x}_{k}(t)\rangle_{\mathcal{Z}_{k}^{(0)}}=0$ and therefore $\hat{m}(t)=0$. Therefore the saddle point equations reduce to Eqs. (3.9) and (3.11).
@@ 3.1.2 Self-consistent stochastic process
where the noise $\eta(t)$ is a Gaussian process with zero mean and variance given by
@@ 3.2 Discrete time dynamical system
$\begin{split}C(t+{\rm d}t,t^{\prime})&=C(t,t^{\prime})+{\rm d}t\left[-\mu(t)C(t,t^{\prime})+g^{2}{\rm d}t\sum_{n=0}^{t^{\prime}/{\rm d}t}f(C(t,n{\rm d}t))R(t^{\prime},n{\rm d}t)\right.\\ &\left.+(J_{0}m(t)+h)m(t^{\prime})\right]\ \ \ \ \ \ \ t>t^{\prime}\\ C(t+{\rm d}t,t+{\rm d}t)&=C(t,t)+2{\rm d}t\left[-\mu(t)C(t,t)+g^{2}{\rm d}t\sum_{n=0}^{t/{\rm d}t}f(C(t,n{\rm d}t))R(t,n{\rm d}t)\right]\\ &+2{\rm d}t(J_{0}m(t)+h)m(t)+{\rm d}t^{2}L(t)\\ R(t+{\rm d}t,t^{\prime})&=R(t,t^{\prime})-{\rm d}t\mu(t)R(t,t^{\prime})+\delta_{t,t^{\prime}}\\ \mu(t)&=\hat{\mu}(C(t,t))\\ m(t+{\rm d}t)&=m(t)+{\rm d}t\left[(J_{0}-\mu(t))m(t)+h\right]\\ L(t)&=\mu^{2}(t)C(t,t)+(J_{0}m(t)+h)^{2}-\mu(t)(J_{0}m(t)+h)m(t)\\ &-2g^{2}\mu(t)\sum_{n=0}^{t/{\rm d}t}f(C(t,n{\rm d}t))R(t,n{\rm d}t){+g^{2}f(C(t,t))}\end{split}$ (3.24)
@@ 6.1 DMFT with periodic drive
$\begin{split}\partial_{t}x_{i}^{(0)}(t)&=-\mu(t)x_{i}^{(0)}(t)+H_{i}(t)\\ \partial_{t}x_{i}^{(1)}(t)&=-\mu(t)x_{i}^{(1)}(t)+gr_{i}({\boldsymbol{x}}(t))\>.\end{split}$ (6.4)
@@
${\mathbb{E}}\left[\cos(\omega s+\theta)\cos(\omega s^{\prime}+\theta)\right]=\frac{1}{2}\cos(\omega(s-s^{\prime})),$ (6.9)
@@
$\begin{split}\partial_{t}C^{(1)}(t,t^{\prime})&=-\mu(t)C^{(1)}(t,t^{\prime})+g^{2}\int_{0}^{t^{\prime}}{\rm d}sf(C(t,s))R^{(1)}(t^{\prime},s)\\ \partial_{t}R^{(1)}(t,t^{\prime})&-\mu(t)R^{(1)}(t,t^{\prime})+\delta(t-t^{\prime})\>.\end{split}$ (6.13)
@@
because in the $N\to\infty$ limit, ${\boldsymbol{x}}^{(0)}(t)\cdot{\boldsymbol{x}}^{(1)}(t^{\prime})/N=0$. This is due to the fact that the dynamics of ${\boldsymbol{x}}^{(0)}$ is autonomous (it does not depend on ${\boldsymbol{x}}^{(1)}$ and the dynamics on ${\boldsymbol{x}}^{(1)}$ depends on ${\boldsymbol{x}}^{(0)}$ only through $\boldsymbol{r}$}. When taking the large $N$ limit one can reduce the dynamics of ${\boldsymbol{x}}^{(1)}$ to an effective single site degree of freedom $x^{1}(t)$ that obeys a self-consistent stochastic equation which depends on ${\boldsymbol{x}}^{(0)}$ only through $C(t,t^{\prime})$. Therefore there is no correlation between these two sets of degrees of freedom. The same property was observed in [38].
@@ 7 Maximum Lyapunov exponent
Chaotic attractors can be characterized by their maximal Lyapunov exponent (MLE), which describes how trajectories are sensible to small perturbations in their initial condition [33].
@@
In this section, we study the problem of determining the MLE for the class of models we are focusing on. We develop a formalism that is alternative to the one of [43] and that partially avoids the field theory techniques that becomes more complicated when dealing with several order parameters. Our technique is similar to [21, 42] and it leads to equations that are close to those presented in [43]. In some of the models that we analyze we can fully compute the analytical expression of the MLE and the asymptotic divergence of the distance between two closely initialized trajectories in terms of the control parameters of the dynamical systems. In the cases in which this is not possible we can still access the MLE by fitting accurately the exponential divergence of two closely initialized trajectories that we can determine via DMFT.
@@ 7.1 Definition of the maximal Lyapunov exponent
$|\delta{\boldsymbol{x}}(t)|\sim{\varepsilon}|{\boldsymbol{b}}_{1}-{\boldsymbol{b}}_{2}|\,e^{\Lambda_{0}(t-t_{0})}\ \ \ \ t\gg t_{0}$ (7.2)
@@ 7.2 DMFT analysis
The distance $D(t,t_{0})$ is a self-averaging quantity [4, 42], indeed one can evaluate it directly by decomposing it into the dynamical correlation functions between the two copies of the system and developing a DMFT treatment.
The two copies are initialized at the same initial point, perform the same trajectory up to time $t_{0}$, get perturbed by the addition of two infinitesimally small random vectors, and then are let free to evolve.
@@ 7.2.1 Two replicas formalism
$x_{\alpha}(t_{0}+{\rm d}t)=x_{\alpha}(t_{0})+{\rm d}t\left[-\mu_{\alpha}(t)x_{\alpha}(t)+\eta_{\alpha}(t)+J_{0}m_{\alpha}(t)+h\right]+{\varepsilon}b_{\alpha},$ (7.10)
@@
$\begin{split}H_{\alpha}(t)&=\begin{cases}{\varepsilon}b_{\alpha}&t=t_{0}\\ 0&t\neq t_{0}\>.\end{cases}\end{split}$ (7.12)
@@
$\hat{H}_{\alpha}(t)={\varepsilon}b_{\alpha}\delta(t-t_{0})$ (7.13)
@@ 7.2.3 Symmetric two replicas DMFT equations
$C_{d}(t+{\rm d}t,t^{\prime})$ $=C_{d}(t,t^{\prime})+{\rm d}t\left[-\tilde{\mu}(t)C_{d}(t,t^{\prime})+g^{2}{\rm d}t\sum_{n=0}^{t^{\prime}/{\rm d}t}f(C_{d}(t,n{\rm d}t))R_{d}(t^{\prime},n{\rm d}t)\right$. (7.22)
@@
$C_{o}(t+{\rm d}t,t^{\prime})$ $=C_{o}(t,t^{\prime})+{\rm d}t\left[-\tilde{\mu}(t)C_{o}(t,t^{\prime})+g^{2}{\rm d}t\sum_{n=0}^{t^{\prime}/{\rm d}t}f(C_{o}(t,n{\rm d}t))R_{d}(t^{\prime},n{\rm d}t)\right$. (7.23)
@@
$C_{d}(t+{\rm d}t,t+{\rm d}t)$ $=C_{d}(t,t)+2{\rm d}t\left[-\tilde{\mu}(t)C_{d}(t,t)+g^{2}{\rm d}t\sum_{n=0}^{t/{\rm d}t}f(C_{d}(t,n{\rm d}t))R_{d}(t,n{\rm d}t)\right]$ (7.24)
@@
$C_{o}(t+{\rm d}t,t+{\rm d}t)$ $=C_{o}(t,t)+2{\rm d}t\left[-\tilde{\mu}(t)C_{o}(t,t)+g^{2}{\rm d}t\sum_{n=0}^{t/{\rm d}t}f(C_{o}(t,n{\rm d}t))R_{d}(t,n{\rm d}t)\right]$ (7.25)